OpenStudy (anonymous):

Two cards are drawn from a standard deck of 52 cards. Find the probability: P(both face cards OR both red) So far I have: 12 total face cards, 26 total red cards, 6 red face cards, C(52,2)=1326. (12*11 + 26*25 - 6)/1326 I checked the answer manual and it should be 188/663, but my work takes me to 388/663. What did I do wrong?

OpenStudy (anonymous):

you counted order. That's what you did wrong

OpenStudy (anonymous):

and it isn't 6 but (6C2)

OpenStudy (anonymous):

WRONG

OpenStudy (anonymous):

Would doing combinations change it? I think I tried earlier, but I can do it again to make sure. C(12,2)= 12!/10!2! = 12*11/2 = 6*11 = 66 C(26,2)= 26!/24!2! = 26*25/2 = 13*25 = 325 C(6,2)= 6!/4!2! = 6*5/2 = 3*5 = 15 Well that's it already. I make the stupidest mistakes sometimes. Gets me to 376, divide by 2 and I have 188. Thanks :B

OpenStudy (anonymous):

no prob